OpenStudy (anonymous):

Explain why the inequalities Nr2 sin(∆φ) ≤ Cr ≤ Nr2 tan(∆φ)

OpenStudy (anonymous):

What does this inequality pertain to?

OpenStudy (anonymous):

here is the entire question

OpenStudy (anonymous):

it is the last one number 5 please look at it and see if you can find anything

OpenStudy (anonymous):

Exactly which part of this is giving you trouble?

OpenStudy (anonymous):

everything about number #5 the last problem with the triangle i can't figure it out at all

OpenStudy (anonymous):

I will get you started, then. The coordinates for Q' are \[(r,r\tan(\theta))\] and the coordinates for Q are \[(r\cos(\theta), r\sin(\theta))\]

OpenStudy (anonymous):

yea that is definitely a start thank you

OpenStudy (anonymous):

For the part with the inequalities, draw pictures of the circles and think about the words "inscribe" and "circumscribe" as they pertain to this situation. Google them if you need to. Normally I'd be more forthcoming and less cryptic with these questions but given that it's a mid-term, I really can't.

OpenStudy (anonymous):

ok thank you very much i really appreciate it

OpenStudy (anonymous):

No problem, good luck

OpenStudy (anonymous):

thanks